I need some help please:
Find the slope of any line perpendicular to the line through points (6, 3) and (7, 7). thank you

Answer by dolly(163)

slope of any line through points (6, 3) and (7, 7) = (7-3)/(7-6)
= 4/1
= 4
So slope of any line perpendicular to the line
through points (6, 3) and (7, 7) = -1/4 [as the product of slopes of perpendicular lines is -1]
The answer is -1/4